OpenStudy (anonymous):

Four of the angles of a pentagon measure 85°, 110°, 135°, and 95°. Find the measure of the missing angle. 115° 95° 135° 85°

OpenStudy (anonymous):

The sum of internal angles in a pentagon is 540.

OpenStudy (anonymous):

whats that supose to mean?

OpenStudy (anonymous):

You can find the missing angle by summing four angles and finding what the fifth angle must be to get to 540.

OpenStudy (anonymous):

i dont get it

OpenStudy (hba):

How many sides are there in a pentagon ?

OpenStudy (anonymous):

5

OpenStudy (hba):

@freebands12 ?

OpenStudy (anonymous):

i think 5

OpenStudy (hba):

Yeah so, Sum of interior angles=(n-2)*180 Here n=5 Sum of interior angles=(5-2)*180 Now calculate

OpenStudy (anonymous):

how do i do that?

OpenStudy (hba):

(5-2)*180 Use a calculator.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

540

OpenStudy (hba):

So, Sum of interior angles=540 Let the unknown angle be "x" so, 85°110°+135°+95+x=540 Now solve for x :)

OpenStudy (anonymous):

what are those little things beside the numbers?

OpenStudy (hba):

That is the degree sign.

OpenStudy (hba):

So,what did you get x as ?

OpenStudy (anonymous):

i didnt get nothing for x i couldnt find anything for x

OpenStudy (hba):

85°+110°+135°+95+x=540 Did you solve for x ?

OpenStudy (hba):

@freebands12 Use a calculator ?

OpenStudy (anonymous):

what do i have to do to find x?

OpenStudy (hba):

add 85+110+135+95 first.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

425

OpenStudy (hba):

425+x=540 Now subtract both sides by 425

OpenStudy (anonymous):

115

OpenStudy (anonymous):

thanks

OpenStudy (hba):

Welcome :)